OpenStudy (exoexoexo):

if m6 = 3/y and m5 = y2/6 then which of the following expresses m in terms of y? (m6= m to the 6th power, m5= m to the 5th power, and y2 = y to the 2nd power)

OpenStudy (anonymous):

\[m ^{6} = 3/y\]\[m ^{6}y = 3\] \[y = 3/ m ^{6}\]

OpenStudy (anonymous):

^^that's for the 1st equation

OpenStudy (exoexoexo):

thank you

OpenStudy (anonymous):

yeah, the second equation should be solved the same way. just get y alone.